4. If Kc for the dissolving of lead chloride is 1.7 x 10-5 at 20oC what are the equilibrium concentrations of lead and chloride ions in solution? Assume there are no ions in solution initially.

                              ????2(?)↔??2+(??)+2??(??).

b) If the initial concentration of ??(??) was NOT zero, would you expect the equilibrium concentration of ??2+(??)to be greater than or less than the concentration found above? Explain your reasoningusing Le Chatlier’s Principle.
Answer #1

a)

At equilibrium:
PbCl2     <---->     Pb2+     +         2 Cl-   


                     s                  2s      


Ksp = [Pb2+][Cl-]^2
1.7*10^-5=(s)*(2s)^2
1.7*10^-5= 4(s)^3
s = 1.62*10^-2 M

SO,
[Pb2+] = s = 1.6*10^-2 M
[Cl-] = 2s
= 2*1.6*10^-2 M
= 3.2*10^-2 M

Answer:
[Pb2+] = 1.6*10^-2 M
[Cl-] = 3.2*10^-2 M

b)
Since there are initial Cl-, which is a product, equilibrium will shift to left according to Le Chatellier’s principle.

So, the concentration of Pb2+ will be less